OpenStudy (anonymous):

help? a man drives his car at a uniform rate from his home toward albany ny. the man;s son left the same house 45 minutes later than his father and traveled along the same road at a rate 12miles per hour faster than his father. the son overtook his father 130 miles from home. what was the rate at which the father drove?

OpenStudy (anonymous):

lool. I'm doing the same type of question. I'm stuck too

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@riccyscaduto

OpenStudy (anonymous):

anyone know what to do? lol

OpenStudy (anonymous):

kinza

OpenStudy (anonymous):

I should be able to do this eventually, but I'm not your best bet for the right answer. :*

OpenStudy (anonymous):

how do i tag someone?

OpenStudy (anonymous):

well i know the answer is either 40 or 2.4 haha

OpenStudy (anonymous):

@ and then username

OpenStudy (anonymous):

its multiple choice , but i would like to know how to actually do the problem

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

I have an way of getting 40, but it's a little messy

OpenStudy (anonymous):

please explain =)

OpenStudy (anonymous):

Give me two minutes to see if there's a neater way of doing it

OpenStudy (anonymous):

ok, take your time

OpenStudy (anonymous):

Got it. :D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

If you let x be the speed that the father is travelling in miles per hour and t be the time travelled in miles/hour, the equation for the distance the father travels is: \[xt\] and for the son it's: \[(x+12)(t-\frac{ 3 }{ 4 })\]

OpenStudy (anonymous):

Because the son travels 12 miles/hour faster than his father, and he takes 45 minutes less to reach that same distance. You have to change 45 minutes to hours, because the speeds are in miles/hour.

OpenStudy (anonymous):

Are you following, Mr. Noodles?

OpenStudy (anonymous):

I'll continue when you are. qq

OpenStudy (anonymous):

ok, i get this so far

OpenStudy (anonymous):

Okay, so those are equations for the distances travelled by both father and son, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We know that the son overtakes his father 130 miles from home--after a distance of 130 miles. So let both equations equal 130.\[xt = 130\]\[(x+12)(t-\frac{ 3 }{ 4 })=130\] and now you have a system of equations you can use to solve for x, which we initially established as the speed the father was travelling.

OpenStudy (anonymous):

once u solve for x?

OpenStudy (anonymous):

wont it be in terms of t?

OpenStudy (anonymous):

=(

OpenStudy (anonymous):

You're solving for t first to find x, right? Because in the equations we've created here, the time taken should be equal--we've taken the 45 minute delay of the son into consideration numerically.

OpenStudy (anonymous):

so u would set xt=(x+12)(t-3/4) equal to eachother?

OpenStudy (anonymous):

Not how I did it. Solve for one variable first--you can't work with two.

OpenStudy (anonymous):

Solve for t first.

OpenStudy (anonymous):

t=130/x

OpenStudy (anonymous):

then plug it in?

OpenStudy (anonymous):

Oh, that could work, yeah.

OpenStudy (anonymous):

is that how you did it?

OpenStudy (anonymous):

Not exactly, but that method works just fine. =)

OpenStudy (anonymous):

Once you end up with a quadratic, tell me what it is and I can confirm it.

OpenStudy (anonymous):

so plugging it in would get you x(130/x)=130?

OpenStudy (anonymous):

or am i not doing this right?

OpenStudy (anonymous):

Haha, you're plugging it into the wrong equation. Why not solve it my way? Expand both equations and isolate for t, then let t=t and solve for x.

OpenStudy (anonymous):

ok that seems easier lol

OpenStudy (anonymous):

so 130/x=(130/x+12)+3/4

OpenStudy (anonymous):

Hm, not quite.

OpenStudy (anonymous):

\[xt = 130\]\[t = \frac{ 130 }{ x }\] \[(x+12)(t-\frac{ 3 }{ 4 })=130\]\[xt -\frac{ 3x }{ 4 }+12t -9=130\]\[(x+12)t-\frac{ 3x }{ 4 }=139\]\[t=\frac{ 139+\frac{ 3x }{ 4 } }{ x+12 }\]

OpenStudy (anonymous):

ok so then what would u do to make it a quadratic?

OpenStudy (anonymous):

Now we know what t is equal to...so t = t and rearrange to create a quadratic to solve for x.

OpenStudy (anonymous):

i did not get an even answer, how do u rearange it to a quadratic? i did not do it correctly

OpenStudy (anonymous):

\[t=t\]\[\frac{ 130 }{ x }=\frac{ 139+\frac{ 3x }{ 4 } }{ x+12 }\]\[(x+12)(130)=(139+\frac{ 3x }{ 4 })(x)\]\[130x+1560=139x+\frac{ 3x^{2} }{ 4 }\]\[0=\frac{ 3x^{2} }{ 4 }+9x-1560\] Use the quadratic equation to solve for x.

OpenStudy (anonymous):

oh ok, so u get -52 and 40 nd have to reject the negative?

OpenStudy (anonymous):

thank you so much!!! how do i give u medals u deserve like 100 lol ?

OpenStudy (anonymous):

Hahahaha, you're very welcome. And of course, his speed can't be negative. Speed isn't a vector, it has no direction, so it's always positive.

OpenStudy (anonymous):

thank you again, finally i can sleep! have a good night =)